OpenStudy (anonymous):

Why doesn't anyone help me? ;( is it because of my name How do you rationalize 3 over 3+ radical 7 - 10 over radical 7!! help help help :)

OpenStudy (anonymous):

i know an aisha and she is very nice

OpenStudy (anonymous):

I am also very nice :)

OpenStudy (anonymous):

I feel like no one helps me because of my name. I can change it if anyone likes?

OpenStudy (anonymous):

\[\frac{3}{3+\sqrt{7}}=\frac{3}{3+\sqrt{7}}\times\frac{3-\sqrt7}{3-\sqrt7}\] \[=\frac{3(3-\sqrt7)}{9-7}=\frac{9-3\sqrt{7}}{2}\]

OpenStudy (anonymous):

Oh you are the best!!!!

OpenStudy (anonymous):

I really appreciate it :) If I ACE that final I owe you one lol

OpenStudy (anonymous):

you rationalize the denominator by multiplying top and bottom by the conjugate of the denominator the conjugate of \(a+\sqrt{b}\) is \(a-\sqrt{b}\) and this works because \[(a+\sqrt b)(a-\sqrt b)=a^2-b\] so there is no more radical in the denominator

OpenStudy (anonymous):

oh okay I got that part but how do you rationalize \[\sqrt{7}\]

OpenStudy (anonymous):

for \[\frac{10}{\sqrt{7}}\] it is easier, just multiply top and bottom by \(\sqrt{7}\) and get \[\frac{10\sqrt{7}}{7}\]

OpenStudy (anonymous):

now your problem is \[\frac{9-3\sqrt{7}}{2}-\frac{10\sqrt{7}}{7}\] so you need a common denominator, namely \(14\)

OpenStudy (anonymous):

OH awesome! so you are left with 9-3\[\sqrt{7}\]/2 - 10\[\sqrt{7}\]/7

OpenStudy (anonymous):

\[\frac{7(9-3\sqrt{7})-2(10\sqrt{7})}{14}\] is the first step

OpenStudy (anonymous):

then multiply out in the numerator and combine like terms

OpenStudy (anonymous):

Oh okay.. wow you're great!

OpenStudy (anonymous):

\[\frac{63-21\sqrt{7}-20\sqrt{7}}{14}=\frac{63-41\sqrt{7}}{14}\]

OpenStudy (anonymous):

hope all steps are clear

OpenStudy (anonymous):

Yes, indeed they are!

OpenStudy (anonymous):

ya but I guess that's not the case anymore after satelite helped me